A charge which has a magnitude of 18 nanoCoulomb produces an electric field. At a distance r the strength of Electric field is equal to . How far is the source charge from the charge?
1
Expert's answer
2020-06-03T12:04:00-0400

The strength of electric field "E" produced by the charge "q" at a distance "r" is given by the equation:


"E = \dfrac{1}{4\pi\varepsilon_0}\dfrac{q}{r^2}"

where "\varepsilon_0 = 8.85\cdot 10^{-12} F/m" is the permittivity of free space.

Thus, for the given charge and strength of electric field the distance will be:


"r = \sqrt{\dfrac{q}{4\pi\varepsilon_0E}}"

Answer. "r = \sqrt{\dfrac{q}{4\pi\varepsilon_0E}}."